The table below shows the results of a survey in which 141 men and 145 women workers ages 25 to 64 were asked # thoy have at least one monthsIncome set anide for omergencies Complete parts (a) through ( Men Women Total 149 137 41 145 286 ess than one months658 income One months income or more Total 76 61 (a) Find the probability that a randomly selected worker has one months income or more set aside for emergencies The probability is 479 (Round to the nearest thousandth as needed) (b) Given that a randomly selected worker is a male, find the probablity hat the worker has less than one months income The probability is Round to the nearest thounandth as needed
Answer #1

(a) Probability = 137/286 = 0.479

(b) Probability = (65/286) / (141/286) = 65/141 = 0.461